How many 1 inch equiltateral triangles are needed to build a chain of triangles that has a perimeter of 20 inches?
Answer by robertb(5830)

x+(x+1)+2 = 20
2x + 3 = 20, or 2x = 17, no integer solution.
x+x+2 = 20,
2x = 18, or x = 9. Therefore you need 2*9 = 18 1 inch equilateral triangles.(9 at the top, 9 at the bottom.)